OpenStudy (anonymous):

Simplify: square root of 108 - 3 square root of 50 + 5 square root of 12 + 4 square root of 18

OpenStudy (valpey):

Try out the equation editor, it is terrific. Do you mean: \[\sqrt{108} + 3\sqrt{50} + 5\sqrt{12} + 4\sqrt{18}\]?

OpenStudy (anonymous):